OpenStudy (anonymous):

please help medal Part B: How many hours did each student learn French each month? Show your work and justify each step of your work. (5 points)

OpenStudy (anonymous):

@butterflydreamer can u please help

OpenStudy (butterflydreamer):

are you missing some information?

OpenStudy (anonymous):

yeah hold on Veronica teaches French in groups of four students. She rents a room for $200 per month. She charges $45 per hour per student. Veronica teaches seven groups of students, and her net earnings per month are $6,100. Part A: If each student learns French for the same number of hours, write an equation that can be used to calculate the number of hours each student learns French in a month. (5 points) Part B: How many hours did each student learn French each month? Show your work and justify each step of your work. (5 points)

OpenStudy (butterflydreamer):

did you work out Part A already?

OpenStudy (anonymous):

yeah can u tell me if Im right 45x - 200 = 6100

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

Salut! Je regarde tu avoir difficulté résolution cette problème. Je ne peut pas aider tu, pour je suis seulement un étudiant d'algèbre, mais je souhaiter tu bonne chance.

OpenStudy (anonymous):

?????

OpenStudy (butterflydreamer):

okay firstly. The question tells us that Veronica teaches 7 GROUPS and each group has 4 STUDENTS. So this means in total she teaches 7 times 4= 28 students. And each hour is $45. So if each student learns french for the same hrs we say let the hrs be x so HER NET INCOME = $ 6100 and she has 28 students who pay $45 per x So 6100= 28 TIMES 45x -200

OpenStudy (anonymous):

so it would be more like 6100=23*45x-200 for part A

OpenStudy (butterflydreamer):

28 ,but i think so.

OpenStudy (anonymous):

where would the 28 go?

OpenStudy (butterflydreamer):

instead of 23

OpenStudy (anonymous):

where does it say that each group has 4 students?

OpenStudy (butterflydreamer):

the first line of teh question

OpenStudy (anonymous):

ok thank u so it would be like this 6100=28*45x-200

OpenStudy (anonymous):

ahh so it does!

OpenStudy (anonymous):

thank you can u help with part B it look right to me

OpenStudy (butterflydreamer):

i think Part B just wants you to solve our equation in part A. so solve for x?

OpenStudy (anonymous):

je blague! je suis très désolé pour perturbateurs votre conversation, bon jour maintenant!

OpenStudy (butterflydreamer):